For an electric circuit with L=0.05 henry, R=20 ohms and C=100*10^-6 farad, the applied emf is 100 volts. Prove that the charge q at time 't' is given by q(t) =0.01-e^(-200t) [0.01 cos(400t) +0.02 sin(400t) ] if initially q=0 and i=0.

1
Expert's answer
2022-03-28T10:59:10-0400

Solution.

Due to Kirchhoff's low:

"U_L+U_R+U_C=E" (1)

where

"U_L=L\frac{di}{dt}=L\frac{d^2q}{dt^2}" ; "U_R=iR=\frac{dq}{dt}R" ; "U_C=\frac1Cq" .

"L\frac{d^2q}{dt^2}+R\frac{dq}{dt}+\frac 1C q=E"

"\frac{d^2q}{dt^2}+\frac RL\frac{dq}{dt}+\frac {1}{LC} q=\frac EL"

"\frac RL=\frac {20}{0.05}=400""\frac {1}{LC}=\frac {1}{0.05\cdot 100\cdot 10^{-6}}=2\cdot10^5"; "\frac EL=\frac{100}{0.05}=2000"

"\frac{d^2q}{dt^2}+400\frac{dq}{dt}+2\cdot10^5q=2000" (2)

"\frac{dq}{dt}=" "\frac{d}{dt}(0.01-e^{-200t} [0.01 \cos(400t) +0.02 \sin(400t) ])" "=e^{-200 t} (-6 \cos(400 t) + 8 \sin(400 t))";

"\frac{d^2q}{dt^2}=400 e^{-200 t} (11 \cos(400 t) + 2 \sin(400 t))".

Substitution all these into (2) gives:

"400 e^{-200 t} (11 \cos(400 t) + 2 \sin(400 t))+" "400\cdot e^{-200 t} (-6 \cos(400 t) + 8 \sin(400 t))+2\cdot 10^5 \cdot" "(0.01-e^{-200t} [0.01 \cos(400t) +0.02 \sin(400t) ])=" "2000"

We see that the left side is equal to the right one, which means "q(t)" satisfies Kirchhoff's law.